In a three-phase, four-wire system, the currents in lines A, B and C under abnormal conditions of loading were as follows:

 Ia = 100∠ 30°  A;            Ib=50∠ 300° A ;    and       Ic=30∠ 180 ° A.

 Determine the zero-phase sequence current  (Iao) in line A in polar form.  


1
Expert's answer
2021-06-07T05:46:46-0400

Write the components matrix

"\begin{bmatrix}
 I_{a_{o} } \\
 I_{a_{1} } \\
 I_{a_{2} }
\end{bmatrix}=\begin{bmatrix}
 1 &1&1\\
 1&k&k^2 \\
 1&k^2&k
\end{bmatrix}\begin{bmatrix}
 I_A \\
 I_B \\
 I_C
\end{bmatrix}"

"k=1\angle120^o; k^2=1\angle120^o"

"I_A= I_{a_{o} } + I_{a_{1} } + I_{a_{2} }"

"I_{a_{o} }=\frac{1}{3} [I_a + I_b+I_c]= 81.777\angle4.693=\frac{81.777}{3}=27.259\angle4.693^o"

"I_{a_{o} }=\frac{1}{3} [I_A +k I_B+k^2I_C]=\frac{1}{3} [100\angle30^o+\angle120^o+50\angle300^o+30\angle180^o]=\frac{173.944\angle43.29^oA}{3} =57.98\angle43.29^oA"

"I_{a_{2} }=\frac{1}{3} [I_A +k^2 I_B+kI_C]=\frac{1}{3} [100\angle30^o+\angle-120^o+50\angle300^o+\angle120^o +30\angle180^o]=\frac{56.929\angle24.96^oA} =18.973\angle24.96"